# The cone conjecture for primitive symplectic varieties over a field of characteristic zero and an application

### v1

@@ 3.1. Preliminaries
First, for $L=\mathbb{C}$, it is known thanks to [undefat, Theorem 3.10]. Next, assume that $L\subset\mathbb{C}$ and denote by $p:X_{\mathbb{C}}\to X$ the projection induced by the base change $X_{\mathbb{C}}$ of $X$ along $\mathop{\rm Spec}\nolimits\mathbb{C}\to\mathop{\rm Spec}\nolimits L$. To simplify notation, denote by $p^{*}:\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits}\xrightarrow{\simeq}\mathop{\rm Pic}\nolimits(X_{\mathbb{C}})_{\mathop{\rm tf}\nolimits}$ the isometry of Lemma 3.9. Note that $X_{\mathbb{C}}$ has terminal singularities thanks to Lemma 3.8. Furthemore, if $E$ is a prime exceptional divisor, then its base change $E_{\mathbb{C}}=p^{-1}(E)$ is still a prime exceptional divisor. Now $r_{E}$ is just $(p^{*})^{-1}\circ r_{E_{\mathbb{C}}}\circ p^{*}$, and since $r_{E_{\mathbb{C}}}$ is an isometry of $\mathop{\rm Pic}\nolimits(X_{\mathbb{C}})_{\mathop{\rm tf}\nolimits}$, it is clear that $r_{E}\in O(\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits})$.
Finally, for the general case, we choose a finitely generated subfield $L^{\prime}\subset L$ over $\mathbb{Q}$ and two primitive symplectic varieties $X^{\prime}$ and $E^{\prime}$ over $\overline{L^{\prime}}\subset L$ such that $X=X^{\prime}_{L}$ and $E=E^{\prime}_{L}$ thanks to Lemma 2.3 and Corollary 2.6. After fixing an embedding $\overline{L^{\prime}}\subset\mathbb{C}$, we know from the previous case that $r_{E^{\prime}}\in O(\mathop{\rm Pic}\nolimits(X^{\prime})_{\mathop{\rm tf}\nolimits})$. Now the same reasoning as above shows that $r_{E}$ is conjugated to $r_{E^{\prime}}$ by the isometry $\pi^{*}:\mathop{\rm Pic}\nolimits(X^{\prime})_{\mathop{\rm tf}\nolimits}\xrightarrow{\simeq}\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits}$ where $\pi:X\to X^{\prime}$ is the projection, therefore $r_{E}\in O(\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits})$ and this completes the proof. ∎
@@
Take a finitely generated subfield $L^{\prime}\subset L$ over $\mathbb{Q}$ such that both $X$ and $E$ are defined over $\overline{L^{\prime}}$, i.e. $X=X^{\prime}_{L}$ and $E=E^{\prime}_{L}$ for $X^{\prime}$ a projective primitive symplectic variety with terminal singularities (Lemma 3.8) over $\overline{L^{\prime}}$ and $E^{\prime}\subset X^{\prime}$ a prime exceptional divisor on $X^{\prime}$. By fixing an embedding $\overline{L^{\prime}}\subset\mathbb{C}$ and considering the projective primitive symplectic variety $X^{\prime}_{\mathbb{C}}$ with the prime exceptional divisor $E^{\prime}_{\mathbb{C}}$ on it, we can see that it is sufficient to deal with the case $L=\mathbb{C}$ since $\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits}$ and $\mathop{\rm Pic}\nolimits(X^{\prime}_{\mathbb{C}})_{\mathop{\rm tf}\nolimits}$ are isometric by Lemma 3.9. We therefore assume that $L=\mathbb{C}$. We will prove something more general by replacing $\mathop{\rm Pic}\nolimits(X)$ with $H^{2}(X,\mathbb{Z})$ in the statement of this lemma. Following the notations of [undefat, Remark 2.9], for $\alpha\in H^{2}(X,\mathbb{Q})$, we write $\alpha^{\vee}\in H_{2}(X,\mathbb{Q})$ for the unique class satisfying $\alpha^{\vee}\cdot\beta=q_{X}(\alpha,\beta)$ for any $\beta\in H^{2}(X,\mathbb{Q})$. In this way, we obtain an isomorphism $H^{2}(X,\mathbb{Q})\to H_{2}(X,\mathbb{Q}),\alpha\mapsto\alpha^{\vee}$. Denote by $[E]\in H^{2}(X,\mathbb{Q})$ the class of $E$. By [undefar, Theorem 1.2], $E$ is uniruled and the dual $[E]^{\vee}$ is proportional to the class of a general curve $R$ in a distinguished ruling. It is also noted in the proof of this theorem that $[E]\cdot[R]=-2$, so if $[E]^{\vee}=c[R]$ with $c\in\mathbb{Q}$, we deduce that
@@
Let $L$ be an algebraically closed field of characteristic $0$ and let $X$ be a primitive symplectic variety over $L$. The subgroup of $O(\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits})$ generated by the reflections with respect to prime exceptional divisors will be denoted by $W_{X}$.
@@
For $L=\mathbb{C}$, the group $W_{X}$ is not defined in this way in [undefat, Lemma 5.4]. Indeed, the authors define it as the subgroup of $\operatorname{Mon}^{2,\operatorname{lt}}_{\operatorname{Hdg}}(X)\subset O(H^{2}(X,\mathbb{Z}),q_{X})$ generated by reflections by classes of stably exceptional line bundles, see [undefat, Definition 2.16] for the notations. However, thanks to [undefat, Theorem 5.12.(3), Theorem 5.12.(5), Lemma 6.3.(2)], the image in $O(\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits})$ of their group is isomorphic to our group, so their definition coincides in the Picard lattice of $X$ with the definition of $W_{X}$ given above.
@@
Let $L$ be an algebraically closed field of characteristic $0$ and let $X$ be a projective primitive symplectic variety over $L$ with $b_{2}(X)\geq 5$ and $\mathbb{Q}$-factorial terminal singularities. Then $\mathop{\overline{\rm Mov}}\nolimits(X)\cap\mathop{\rm Pos}\nolimits(X)$ is a fundamental domain for the action of $W_{X}$ on $\mathop{\rm Pos}\nolimits(X)$, cut out by closed half-spaces associated to prime exceptional divisors.
@@
For the case $L=\mathbb{C}$, the result is known thanks to [undefat, Lemma 6.2]. For the general case, take a finitely generated subfield $L^{\prime}\subset L$ over $\mathbb{Q}$ and a $X^{\prime}$ a primitive symplectic variety over $L^{\prime}$ such that $X=X^{\prime}_{L}$. After fixing an embedding $\overline{L^{\prime}}\subset\mathbb{C}$, Lemma 3.9 gives isometries
$\mathop{\rm Pic}\nolimits(X)_{\mathbb{R}}\simeq\mathop{\rm Pic}\nolimits\left(X^{\prime}_{\overline{L^{\prime}}}\right)_{\mathbb{R}}\simeq\mathop{\rm Pic}\nolimits\left(X^{\prime}_{\mathbb{C}}\right)_{\mathbb{R}}$
which are induced by the pullbacks along the projections of the base changes. Observe that $X^{\prime}$ and $X^{\prime}_{\mathbb{C}}$ have $\mathbb{Q}$-factorial and terminal singularities thanks to Lemma 3.8, and that $b_{2}(X^{\prime}),b_{2}(X^{\prime}_{\mathbb{C}})\geq 5$ by [undefaac, VI - Corollary 4.3]. Given that the movable cone (respectively the positive cone) of $X^{\prime}_{\mathbb{C}}$ corresponds bijectively to that of $X^{\prime}_{\overline{L^{\prime}}}$ and $X$ by the above isometries thanks to Lemma 3.4, the result follows from the case $L=\mathbb{C}$. ∎
Definition 3.15.
Let $F$ be a field of characteristic $0$ and let $X$ be a primitive symplectic variety over $F$ with terminal singularities. Let $\overline{X}:=X_{\overline{F}}$. Consider $W_{\overline{X}}\subset O\left(\mathop{\rm Pic}\nolimits\left(\overline{X}\right)_{\mathop{\rm tf}\nolimits}\right)$ and denote by $G_{F}$ the absolute Galois group $\mathop{\rm Gal}\nolimits\left(\overline{F}/F\right)$ of $F$. Then $G_{F}$ acts via automorphisms on $\mathop{\rm Pic}\nolimits\left(\overline{X}\right)$ inducing a group morphism $\alpha:G_{F}\to O\left(\mathop{\rm Pic}\nolimits\left(\overline{X}\right)_{\mathop{\rm tf}\nolimits}\right)$. Thanks to this morphism, $G_{F}$ acts by conjugation on $O\left(\mathop{\rm Pic}\nolimits\left(\overline{X}\right)_{\mathop{\rm tf}\nolimits}\right)$:
@@
As explained in [undefm], it is clear that $R_{X}$ induces an action on $\mathop{\rm Pic}\nolimits\left(\overline{X}\right)_{\mathop{\rm tf}\nolimits}^{G_{F}}$ but it is not entirely clear that this action preserves $\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits}$. To prove this, we would like to apply the arguments used to prove [undefm, Proposition 3.6]. We will need the notion of a Coxeter system which we recall below.
@@
Let $W$ be a group with identity element $1$ and let $T\subset W$ be a set which generates $W$, such that $T=T^{-1}$ and $1\notin T$. For any $w\in W$, the length $l(w)$ of $w$ is the smallest integer $q\geq 0$ such that $w$ is a product of a sequence of $q$ elements of $T$. Assume that all the elements of $T$ have order $2$. For $t_{i},t_{j}\in T$, let $n_{i,j}=n_{j,i}$ be the order of $t_{i}t_{j}$ if it is finite, $0$ otherwise. We say that $(W,T)$ is a Coxeter system if the generating set $T$ and the relations $(t_{i}t_{j})^{n_{i,j}}=1$ for $n_{i,j}\neq 0$ form a presentation of the group $W$.
@@
The Coxeter–Dynkin diagram of a Coxeter system $(W,T)$ is the graph obtained by taking the set $T$ as the set of vertices, and declaring that two vertices $t_{i}$ and $t_{j}$ are adjacent if and only if the commutator $[t_{i},t_{j}]$ in $W$ is not equal to $1$. Each edge between $t_{i}$ and $t_{j}$ is weighted by $n_{i,j}-2$ if $n_{i,j}>0$, and by $0$ otherwise.
@@
    If $W$ is a finite group, there exists a unique $w_{0}\in W$ such that for any $w\in W\backslash\{w_{0}\}$, we have $l(w)<l(w_{0})$. We say that $w_{0}$ is the longest element of $(W,T)$. [undefg, Proposition 2.3.1]
@@
    If $I$ is a non-empty subset of $T$, let $W_{I}$ denote the subgroup of $W$ generated by the elements of $I$. Then $(W_{I},I)$ is a Coxeter system. [undefg, Proposition 2.4.1.(i)]
The strategy is then to adapt [undefaat, Proposition 4.1.3], relying on [undefm, Proposition 3.6], to the singular case. In the next proposition, we will make the slight abuse of notation of using the same letter to denote a set consisting of exceptional prime divisors and the set of associated reflections.
@@
Let $F$ be a field of characteristic $0$ and let $X$ be a primitive symplectic variety over $F$ with $\mathbb{Q}$-factorial and terminal singularities. Denote by $G_{F}$ the absolute Galois group of $F$. Let $I$ be a Galois orbit of prime exceptional divisors on $\overline{X}:=X_{\overline{F}}$ and denote by $W_{I}$ the subgroup of $W_{\overline{X}}$ generated by the reflections with respect to elements of $I$. Let $\Phi$ be the set of Galois orbits $I$ of prime exceptional divisors such that $W_{I}$ finite. The following holds.
@@
    If ${\mathcal{P}}_{\overline{X}}$ is the set of prime exceptional divisors of $\overline{X}$, then both $(W_{\overline{X}},{\mathcal{P}}_{\overline{X}})$ and $(W_{I},I)$ are Coxeter systems.
@@
    For each $I\in\Phi$, let $r_{I}$ be the longest element of the Coxeter system $(W_{I},I)$ as in Proposition 3.18, and let $E_{I}$ be the sum of the elements in $I$. Then $\left(R_{X},\left\{r_{I}\penalty 10000\ \middle|\penalty 10000\ I\in F\right\}\right)$ is a Coxeter system and $r_{I}$ acts on $\mathop{\rm Pic}\nolimits\left(\overline{X}\right)^{G_{F}}_{\mathop{\rm tf}\nolimits}$ as the reflection in the class of a sufficiently divisible integral multiple of $E_{I}$.
@@
(1) This is a consequence of Proposition 3.18.(2) and [undefah, Section 5.4].
(2) Assume that $W_{I}$ is finite and take two distinct elements $E_{1}$ and $E_{2}$ in $I$. Up to replacing $\overline{F}$ by a finitely generated subfield over which $\overline{X}$, $E_{1}$ and $E_{2}$ are defined, the Lefschetz principle allows us to assume that $\overline{F}=\mathbb{C}$. In this case, the classes of $E_{1}$ and $E_{2}$ in $\mathop{\rm Pic}\nolimits(\overline{X})_{\mathbb{R}}$ are different by [undefal, Theorem 1.1]. To simplify the notation, let $\alpha=q_{\overline{X}}(E_{1},E_{2})$ and $\beta=q_{\overline{X}}(E_{1},E_{1})=q_{\overline{X}}(E_{2},E_{2})$. Reflections $r_{E_{1}}$ and $r_{E_{2}}$ stabilize the $2$-plane $W:=\mathbb{R}E_{1}\oplus\mathbb{R}E_{2}$. After restricting these reflections to $W$, we obtain the matrices
@@
These are matrices with integer coefficients by Lemma 3.11 and, by assumption, they have finite order. By classical results in linear algebra, these orders belong to the set {2,3,4,6}. The order of $r_{E_{1}}r_{E_{2}}$ cannot be equal to either $4$ or $6$, as can be seen by computing the determinant and the trace from the eigenvalues, which are conjugate primitive roots of unity. A calculation shows that it is equal to $2$ if and only if $\alpha=0$, and that it is equal to $3$ if and only if $\beta=-2\alpha$. By [undefg, Exercise 1.4], the Coxeter–Dynkin diagram of $(W_{I},I)$ is a finite union of finite trees, and we can therefore deduce that there is a vertex of degree at most $1$. But now the Galois group $G_{F}$ acts transitively on the diagram, so it follows that each vertex has degree at most $1$. On the one hand, saying that this degree is $0$ is equivalent, by definition of the Coxeter-Dynkin diagram, to the fact that the order of $r_{E_{1}}r_{E_{2}}$ is equal to $2$, i.e. $\alpha=0$ and we are then in situation (2a) of the statement. On the other hand, saying that this degree is $1$ is equivalent to the fact that the order of $r_{E_{1}}r_{E_{2}}$ is equal to $3$, i.e. $\beta=-2\alpha$ and we are then in situation (2b) of the statement.
(3) $\left(R_{X},\left\{r_{I}\penalty 10000\ \middle|\penalty 10000\ I\in F\right\}\right)$ is a Coxeter system by [undefz, Theorem 1]. As for the action of $r_{I}$, we prove it separately for the two types (2a) and (2b) of orbits. In the first case (2a), we have $I=\{E_{1},\dotsc,E_{r}\}$ and $E_{I}=E_{1}+\dotsb+E_{r}$. After replacing $E_{I}$ by a sufficiently divisible multiple, we may assume that $E_{I}$ is Cartier. The group $W_{I}$ is isomorphic to the Coxeter group $A_{1}^{r}$ and the longest element is therefore equal to $r_{I}=r_{E_{1}}\circ\dotsb\circ r_{E_{r}}$. For $\alpha\in\mathop{\rm Pic}\nolimits(\overline{X})^{G_{F}}_{\mathop{\rm tf}\nolimits}$, we obtain

### v2

@@ 3.1. Preliminaries
First, the case $L=\mathbb{C}$ is dealt with in [undefas, Theorem 3.10]. Next, assume that $L\subset\mathbb{C}$ and denote by $p:X_{\mathbb{C}}\to X$ the projection. Note that $X_{\mathbb{C}}$ has terminal singularities thanks to Lemma 3.8. Furthemore, if $E$ is a prime exceptional divisor, then its base change $E_{\mathbb{C}}\subset X_{\mathbb{C}}$ is still a prime exceptional divisor. Let $p^{*}:\mathop{\rm Pic}\nolimits(X)_{\mathbb{Q}}\xrightarrow{\simeq}\mathop{\rm Pic}\nolimits(X_{\mathbb{C}})$ be the induced isometry of Lemma 3.9. Since $r_{E}$ is equal to $(p^{*})^{-1}\circ r_{E_{\mathbb{C}}}\circ p^{*}$ and since $r_{E_{\mathbb{C}}}$ induces by restriction an isometry of $\mathop{\rm Pic}\nolimits(X_{\mathbb{C}})_{\mathop{\rm tf}\nolimits}$, it follows that $r_{E}$ induces an isometry of $\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits}$.
Finally, for the general case, take a finitely generated subfield $L^{\prime}\subset L$ over $\mathbb{Q}$ and two primitive symplectic varieties $X^{\prime}$ and $E^{\prime}$ over $\overline{L^{\prime}}\subset L$ such that $X=X^{\prime}_{L}$ and $E=E^{\prime}_{L}$ thanks to Lemma 2.3 and Corollary 2.6. By fixing an embedding $\overline{L^{\prime}}\subset\mathbb{C}$, it follows that $r_{E^{\prime}}\in O(\mathop{\rm Pic}\nolimits(X^{\prime})_{\mathop{\rm tf}\nolimits})$ by the previous case. Since $r_{E}$ is conjugated as above to $r_{E^{\prime}}$ by the isometry $\pi^{*}:\mathop{\rm Pic}\nolimits(X^{\prime})_{\mathbb{Q}}\xrightarrow{\simeq}\mathop{\rm Pic}\nolimits(X)_{\mathbb{Q}}$ where $\pi:X\to X^{\prime}$ is the projection, we conclude that $r_{E}\in O(\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits})$, and this completes the proof. ∎
@@
Take a finitely generated subfield $L^{\prime}\subset L$ over $\mathbb{Q}$ such that both $X$ and $E$ are defined over $\overline{L^{\prime}}$, i.e. $X=X^{\prime}_{L}$ and $E=E^{\prime}_{L}$ for $X^{\prime}$ a projective primitive symplectic variety with terminal singularities (Lemma 3.8) over $\overline{L^{\prime}}$ and $E^{\prime}\subset X^{\prime}$ a prime exceptional divisor on $X^{\prime}$. By fixing an embedding $\overline{L^{\prime}}\subset\mathbb{C}$ and considering the projective primitive symplectic variety $X^{\prime}_{\mathbb{C}}$ with the prime exceptional divisor $E^{\prime}_{\mathbb{C}}$ on it, we see that it is sufficient to deal with the case $L=\mathbb{C}$ since $\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits}$ and $\mathop{\rm Pic}\nolimits(X^{\prime}_{\mathbb{C}})_{\mathop{\rm tf}\nolimits}$ are isometric by Lemma 3.9. We therefore assume that $L=\mathbb{C}$. We will prove something more general by replacing $\mathop{\rm Pic}\nolimits(X)$ with $H^{2}(X,\mathbb{Z})$ in the statement of this lemma. Following the notations of [undefas, Remark 2.9], for $\alpha\in H^{2}(X,\mathbb{Q})$, we write $\alpha^{\vee}\in H_{2}(X,\mathbb{Q})$ for the unique class satisfying $\alpha^{\vee}\cdot\beta=q_{X}(\alpha,\beta)$ for any $\beta\in H^{2}(X,\mathbb{Q})$. In this way, we obtain an isomorphism $H^{2}(X,\mathbb{Q})\to H_{2}(X,\mathbb{Q}),\alpha\mapsto\alpha^{\vee}$. Denote by $[E]\in H^{2}(X,\mathbb{Q})$ the class of $E$. By [undefaq, Theorem 1.2], $E$ is uniruled and the dual $[E]^{\vee}$ is proportional to the class of a general curve $R$ in a distinguished ruling. It is also noted in the proof of this theorem that $[E]\cdot[R]=-2$, so if $[E]^{\vee}=c[R]$ with $c\in\mathbb{Q}$, we deduce that
@@
Let $L$ be an algebraically closed field of characteristic $0$ and $X$ a projective primitive symplectic variety over $L$. The subgroup of $O(\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits})$ generated by the reflections with respect to prime exceptional divisors will be denoted by $W_{X}$.
@@
For $L=\mathbb{C}$, the group $W_{X}$ is not defined in this way in [undefas, Lemma 5.4]. Indeed, the authors define it as the subgroup of $\operatorname{Mon}^{2,\operatorname{lt}}_{\operatorname{Hdg}}(X)\subset O(H^{2}(X,\mathbb{Z}),q_{X})$ generated by reflections by classes of stably exceptional line bundles, see [undefas, Definition 2.16] for the notations. However, thanks to [undefas, Theorem 5.12.(3), Theorem 5.12.(5), Lemma 6.3.(2)], the image in $O(\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits})$ of their group is isomorphic to our group, so their definition coincides in the Picard lattice of $X$ with the definition of $W_{X}$ given above.
@@
Let $L$ be an algebraically closed field of characteristic $0$ and let $X$ be a projective primitive symplectic variety over $L$ with $b_{2}(X)\geq 5$ and $\mathbb{Q}$-factorial terminal singularities. Then $\mathop{\overline{\rm Mov}}\nolimits(X)\cap\mathop{\rm Pos}\nolimits(X)$ is a fundamental domain under the action of $W_{X}$ on $\mathop{\rm Pos}\nolimits(X)$, cut out by closed half-spaces associated to prime exceptional divisors.
@@
The case $L=\mathbb{C}$ is dealt with in [undefas, Lemma 6.2]. For the general case, take a finitely generated subfield $L^{\prime}\subset L$ over $\mathbb{Q}$ and a $X^{\prime}$ a primitive symplectic variety over $\overline{L^{\prime}}$ such that $X=X^{\prime}_{L}$. After fixing an embedding $\overline{L^{\prime}}\subset\mathbb{C}$, Lemma 3.9 gives isometries
$\mathop{\rm Pic}\nolimits(X)_{\mathbb{R}}\simeq\mathop{\rm Pic}\nolimits\left(X^{\prime}\right)_{\mathbb{R}}\simeq\mathop{\rm Pic}\nolimits\left(X^{\prime}_{\mathbb{C}}\right)_{\mathbb{R}}$
which are induced by the pullbacks along the projections of the base changes. Observe that both $X^{\prime}$ and $X^{\prime}_{\mathbb{C}}$ have $\mathbb{Q}$-factorial and terminal singularities thanks to Lemma 3.8, and that $b_{2}(X^{\prime})=b_{2}(X^{\prime}_{\mathbb{C}})\geq 5$ by [undefaac, VI - Corollary 4.3]. The movable cones (respectively the positive cones) of $X$, $X^{\prime}$ and $X^{\prime}_{\mathbb{C}}$ correspond bijectively through the above isometries by Lemma 3.3, therefore the result follows from the case $L=\mathbb{C}$. ∎
For a projective primitive symplectic variety $X$ with terminal singularities over a field $F$ of characteristic $0$, denote by $G_{F}$ the absolute Galois group of $F$, and by $\overline{X}$ the base change $X_{\overline{F}}$. Let $W_{\overline{X}}\subset O\left(\mathop{\rm Pic}\nolimits(\overline{X})_{\mathop{\rm tf}\nolimits}\right)$ be the subgroup generated by reflections with respect to prime exceptional divisors on $\overline{X}$ as in Definition 3.12. $G_{F}$ acts naturally by automorphisms on $\mathop{\rm Pic}\nolimits(\overline{X})$, and therefore acts naturally by conjugation on $O(\mathop{\rm Pic}\nolimits(\overline{X})_{\mathop{\rm tf}\nolimits})$:
@@
Definition 3.15.
With the above notations, we denote the fixed part of $W_{\overline{F}}$ under this action by
@@
As explained in [undefl], although it is straightforward to see that $R_{X}$ induces an action on $\mathop{\rm Pic}\nolimits\left(\overline{X}\right)_{\mathop{\rm tf}\nolimits}^{G_{F}}$, it is not entirely clear that this action preserves $\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits}$. To prove this, we would like to adapt the arguments used to prove [undefl, Proposition 3.6]. We will need the notion of a Coxeter system which we recall below.
@@
Let $W$ be a group with identity element $1$, and let $T\subset W$ be a generating set such that $T=T^{-1}$ and $1\notin T$. For any $w\in W$, the length $l(w)$ of $w$ is the smallest integer $q\geq 0$ such that $w$ can be written as a product of $q$ elements of $T$. Assume that every element of $T$ has order $2$. For $t_{i},t_{j}\in T$, let $n_{i,j}=n_{j,i}$ denote the order of $t_{i}t_{j}$ if it is finite, and set $n_{i,j}=0$ otherwise. We say that $(W,T)$ is a Coxeter system if $T$ together with the relations $(t_{i}t_{j})^{n_{i,j}}=1$ for $n_{i,j}\neq 0$ gives a presentation of $W$.
@@
The Coxeter–Dynkin diagram of a Coxeter system $(W,T)$ is the graph whose vertex set is $T$, with an edge between two distinct vertices $t_{i}$ and $t_{j}$ if and only if $t_{i}$ and $t_{j}$ do not commute. Each edge joining $t_{i}$ and $t_{j}$ is labelled by $n_{i,j}-2$ if $n_{i,j}>0$, and by $0$ otherwise.
@@
    If $W$ is a finite group, there exists a unique $w_{0}\in W$ such that for any $w\in W\backslash\{w_{0}\}$, we have $l(w)<l(w_{0})$. We say that $w_{0}$ is the longest element of $(W,T)$. [undeff, Proposition 2.3.1]
@@
    If $I$ is a non-empty subset of $T$, let $W_{I}$ denote the subgroup of $W$ generated by the elements of $I$. Then $(W_{I},I)$ is a Coxeter system. [undeff, Proposition 2.4.1.(i)]
The strategy is then to adapt [undefaat, Proposition 4.1.3], relying on [undefl, Proposition 3.6], to the singular case. In the next proposition, we will make the slight abuse of notation of using the same letter to denote a set consisting of exceptional prime divisors and the set of associated reflections.
@@
Let $F$ be a field of characteristic $0$, and let $X$ be a primitive symplectic variety over $F$ with $b_{2}(X)\geq 5$ such that $\overline{X}:=X_{\overline{F}}$ has $\mathbb{Q}$-factorial terminal singularities. Denote by $G_{F}$ the absolute Galois group of $F$. Let $I$ be a Galois orbit of prime exceptional divisors on $\overline{X}$, and denote by $W_{I}$ the subgroup of $W_{\overline{X}}$ generated by the reflections with respect to the elements of $I$. Let $\Phi$ be the set of Galois orbits $I$ of prime exceptional divisors such that $W_{I}$ is finite, and let ${\mathcal{P}}_{\overline{X}}$ be the set of prime exceptional divisors of $\overline{X}$. Then the following holds.
@@
    Both $(W_{\overline{X}},{\mathcal{P}}_{\overline{X}})$ and $(W_{I},I)$ are Coxeter systems.
@@
    For each $I\in\Phi$, let $r_{I}$ be the longest element of the Coxeter system $(W_{I},I)$ as in Proposition 3.18, and let $E_{I}$ be the sum of the elements in $I$. Then $\left(R_{X},\left\{r_{I}\penalty 10000\ \middle|\penalty 10000\ I\in\Phi\right\}\right)$ is a Coxeter system and $r_{I}$ acts on $\mathop{\rm Pic}\nolimits\left(\overline{X}\right)^{G_{F}}_{\mathop{\rm tf}\nolimits}$ as the reflection in the class of a sufficiently divisible integral multiple of $E_{I}$.
@@
(1) This is a consequence of Proposition 3.18.(2) and [undefag, Section 5.4].
(2) Assume that $W_{I}$ is finite and let $E_{1}$ and $E_{2}$ be two distinct elements in $I$. Up to replacing $\overline{F}$ by a finitely generated subfield over which $\overline{X}$, $E_{1}$ and $E_{2}$ are defined, the Lefschetz principle allows us to assume that $\overline{F}=\mathbb{C}$. In this case, the classes of $E_{1}$ and $E_{2}$ in $\mathop{\rm Pic}\nolimits(\overline{X})_{\mathbb{R}}$ are different by [undefak, Theorem 1.1]. To simplify the notation, let $\alpha=q_{\overline{X}}(E_{1},E_{2})$ and $\beta=q_{\overline{X}}(E_{1},E_{1})=q_{\overline{X}}(E_{2},E_{2})$. Reflections $r_{E_{1}}$ and $r_{E_{2}}$ stabilize the $2$-plane $W:=\mathbb{R}E_{1}\oplus\mathbb{R}E_{2}$. Restricting these reflections to $W$, we obtain the matrices
@@
These are matrices with integer coefficients by Lemma 3.11 and, by assumption, they have finite order. By classical results in linear algebra, these orders belong to the set {2,3,4,6}. The order of $r_{E_{1}}r_{E_{2}}$ cannot be equal to either $4$ or $6$, as can be seen by computing the determinant and the trace from the eigenvalues, which are conjugate primitive roots of unity. A calculation shows that it is equal to $2$ if and only if $\alpha=0$, and that it is equal to $3$ if and only if $\beta=-2\alpha$. By [undeff, Exercise 1.4], the Coxeter–Dynkin diagram of $(W_{I},I)$ is a finite union of finite trees, and we can therefore deduce that there is a vertex of degree at most $1$. But now the Galois group $G_{F}$ acts transitively on the diagram, so it follows that each vertex has degree at most $1$. On the one hand, saying that this degree is $0$ is equivalent, by definition of the Coxeter-Dynkin diagram, to the fact that the order of $r_{E_{1}}r_{E_{2}}$ is equal to $2$, i.e. $\alpha=0$ and we are then in situation (2a) of the statement. On the other hand, saying that this degree is $1$ is equivalent to the fact that the order of $r_{E_{1}}r_{E_{2}}$ is equal to $3$, i.e. $\beta=-2\alpha$ and we are then in situation (2b) of the statement.
(3) $\left(R_{X},\left\{r_{I}\penalty 10000\ \middle|\penalty 10000\ I\in\Phi\right\}\right)$ is a Coxeter system by [undefy, Theorem 1]. As under the action of $r_{I}$, we prove it separately for the two types (2a) and (2b) of orbits. In the first case (2a), we have $I=\{E_{1},\dotsc,E_{r}\}$ and $E_{I}=E_{1}+\dotsb+E_{r}$. Replacing $E_{I}$ by a sufficiently divisible multiple, we may assume that $E_{I}$ is Cartier. The group $W_{I}$ is isomorphic to the Coxeter group $A_{1}^{r}$ and the longest element is therefore equal to $r_{I}=r_{E_{1}}\circ\dotsb\circ r_{E_{r}}$. For $\alpha\in\mathop{\rm Pic}\nolimits(\overline{X})^{G_{F}}_{\mathop{\rm tf}\nolimits}$, we obtain